Decide whether the given statement is true or false. If​ false, correct the right side of the equation.
(-5+7i)-(10-6i)=-15+i

Answer by stanbon(75887)

Decide whether the given statement is true or false. If​ false, correct the right side of the equation.
(-5+7i)-(10-6i)=-15+i
-5+7i-10+6i = -15+i
-15+13i = -15+13i
--------------------
Cheers,
Stan H.
---------